# RA-PbRL: Provably Efficient Risk-Aware Preference-Based Reinforcement Learning

### v1

@@ 5 Experiment Results
In this section, we assess the empirical performance of RA-PbRL (Algorithm 1). For a comparative analysis, we select two baseline algorithms: PbOP, as described in Chen et al. (2022), which is a PbRL algorithm utilizing general function approximation, and ICVaR-RLHF, detailed in Chen et al. (2023), which is a risk-sensitive Human Feedback RL algorithm. These baselines represent the most closely aligned algorithms with RA-PbRL, especially in terms of employing general function approximation. The evaluation of empirical performance is conducted through the lens of static regret, as defined in Eq. 8.
@@ 5.2 Experiment settings: Half Cheetah
Aditionally, we implement our algorithm to solve MuJoCo’s Half-cheetah simulation. We implement our proposed algorithm alongside PbOP and TRC. The objective of the algorithms is to learn to control a robot simulation of a cheetah to run forward. This problem differs from the previous setting in that it is more challenging and uses a continuous state-action space. Because the algorithm ws originally implemented for discrete state-action spaces, we assume the transition, reward, and policy functions can be parameterized by a linear function, which is optimized for using gradient descent.
@@ 5.3 Experimental results
As depicted in Figure 1 and 3, the regret of RA-PbRL over static and nested CVaR initially exhibits a linear growth with respect to $K$, transitioning to sublinear growth upon reaching a certain threshold. This behavior aligns with the conclusions drawn in Section 4.2. It is important to note that increased risk aversion ($\alpha\rightarrow 0$) introduces greater uncertainty, as evidenced by the larger variance regions observed in the experiments. Notably, the regret of the bad-scenarios associated with RA-PbRL is significantly lower compared to those of other algorithms. It can additionally be observed that as $\alpha$ is increased, the regret improves, which is expected as riskier behavior can also improve the odds of finding useful behaviors.
@@ 6 Conclusion
In this paper, we investigate a novel PbRL algorithm for solving problems requiring risk awareness. We explore static and nested measures to introduce risk awareness to PbRL settings. To the best of our knowledge, our proposed RA-PbRL algorithm is the first provably efficient Preference-based Reinforcement Learning (PbRL) algorithm that incorporates both nested and static risk objectives in one algorithm. Our algorithm is built on innovative techniques for the efficient approximation of regret. A core finding in our investigation is the strong influence of the state and trajectory dimensions with respect to the nested risk objective regret. On the other hand, the static risk objective regret is mainly determined by the quantile function.
We have also identified the following four limitations to our work. (1) Our comparison feedback is limited to two trajectories. An interesting, more general approach could consider n-wise comparisons. (2) The reward functions are assumed to be linear in this work for the sake of simplicity. (3) Although this work has considered more general risk measures, we have still made certain assumptions that limit the generality of our results. (4) There is still room for future improvements to further close the gap between upper and lower bounds. We believe this work opens several avenues for future research, including establishing the concrete lower bounds of risk-aware PbRL, improving the computational complexity of the algorithm, and conducting experiments in more diverse and interesting environments/simulations.
@@ Appendix A Notation
$\tilde{V}_{h}^{\pi}(s)$ Value function of state $s$ at step $h$ under policy $\pi$ in PbRL-MDP.
@@ Appendix B Risk Aware Object Computability
Unlike standard RL where each step’s reward can be observed, PbRL represents a type of RL characterized by once-per-episode feedback. As a result, our observable and estimable parameters are confined to the trajectory reward $r^{\star}_{\xi}$ and transition probability functions $\textbf{P}^{\star}$. Consequently, the traditional risk-aware objective might be unsuitable, as the reduction in available information prevents the computation of the original risk-aware measure. The risk measure selected for PbRL must satisfy the following condition: it should remain unique across MDPs where policies, trajectory rewards and transition probability functions even when each trajectory is fixed, but different step rewards, $r^{\star}(s,a)$, vary. This requirement ensures that the risk measure consistently reflects the underlying preferences regardless of variations in specific step rewards.
@@ B.1 Nested Object
Initial case. Using the tree-like PbRL-MDP algorithm and the initial conditions of the Bellman equation, at the final step $h=H$, we have
@@ E.1 Regret Lower Bound of Nested Reward
According to the definition of nested-CVaR-PbRL risk objective in Eq. 4, we have:
@@
According to the definition of nested-CVaR-PbRL risk objective in 4, we have:
@@
Appendix F The optimal policy calculation for known PbRL MDP
In this section, we describe how to compute the optimal policy for the CVaR objective when the PbRL-MDP is known; this approach is described in detail in Bäuerle & Ott (2011); Bastani et al. (2022). Following this work, we consider the setting where we are trying to minimize cost rather than maximize reward. In particular, consider an know PbRL MDP $\mathcal{M}(\mathcal{S},\mathcal{A},r_{\xi}^{\star},\mathbf{P}^{\star},H)$, where $\mathcal{S}$ and $\mathcal{A}$, and our goal is to compute a policy $\pi$ that maximizes its CVaR objective.
@@ F.1 Static CVaR object
Value iteration. we reconstruct the MDP as enlarge the state space $\tilde{s}_{h}=(\xi_{h},\rho)$, where $\rho$ will work as a quantile value. Letting $S_{1}$ be the initial state of the original PbRL MDP $\mathcal{M}$
@@ F.2 Nested CVaR object
Value iteration. we reconstruct the MDP as enlarge the state space $\tilde{s}_{h}=(\xi_{h},\rho)$, where $\rho$ will work as a quantile value. Letting $S_{1}$ be the initial state of the original PbRL MDP $\mathcal{M}$
@@ 3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contribution and scope: RA-PbRL algorithm.

### v2

@@ 5 Experiment Results
In this section, we assess the empirical performance of RA-RLHF (Algorithm 1). For a comparative analysis, we select two baseline algorithms: PbOP, as described in Chen et al. (2022), which is a RLHF algorithm utilizing general function approximation, and ICVaR-RLHF, detailed in Chen et al. (2023), which is a risk-sensitive Human Feedback RL algorithm. These baselines represent the most closely aligned algorithms with RA-RLHF, especially in terms of employing general function approximation. The evaluation of empirical performance is conducted through the lens of static regret, as defined in Eq. 8.
@@ 5.2 Experiment settings: Half Cheetah
Additionally, we implement our algorithm to solve MuJoCo’s Half-cheetah simulation. We implement our proposed algorithm alongside PbOP (Preference-based Optimistic Planning,  Chen et al. (2022)) and TRC (Trust-Region CVaR,  Kim & Oh (2022)). The objective of the algorithms is to learn to control a robot simulation of a cheetah to run forward. This problem differs from the previous setting in that it is more challenging and uses a continuous state-action space. Because the algorithm was originally implemented for discrete state-action spaces, we assume the transition, reward, and policy functions can be parameterized by a linear function, which is optimized for using gradient descent.
@@ 5.3 Experimental results
As depicted in Figure 1 and 3, the regret of RA-RLHF over static and nested CVaR initially exhibits a linear growth with respect to $K$, transitioning to sublinear growth upon reaching a certain threshold. This behavior aligns with the conclusions drawn in Section 4.2. It is important to note that increased risk aversion ($\alpha\rightarrow 0$) introduces greater uncertainty, as evidenced by the larger variance regions observed in the experiments. Notably, the regret of the bad-scenarios associated with RA-RLHF is significantly lower compared to those of other algorithms. It can additionally be observed that as $\alpha$ is increased, the regret improves, which is expected as riskier behavior can also improve the odds of finding useful behaviors.
@@ 6 Conclusion
In this paper, we investigate a novel RLHF algorithm for solving problems requiring risk awareness. We explore static and nested measures to introduce risk awareness to RLHF settings. To the best of our knowledge, our proposed RA-RLHF algorithm is the first provably efficient Reinforcement Learning Human Feedback (RLHF) algorithm that incorporates both nested and static risk objectives in one algorithm. Our algorithm is built on innovative techniques for the efficient approximation of regret. A core finding in our investigation is the strong influence of the state and trajectory dimensions with respect to the nested risk objective regret. On the other hand, the static risk objective regret is mainly determined by the quantile function.
We have also identified the following four limitations to our work. (1) Our comparison feedback is limited to two trajectories. An interesting, more general approach could consider n-wise comparisons. (2) The reward functions are assumed to be linear in this work for the sake of simplicity. (3) Although this work has considered more general risk measures, we have still made certain assumptions that limit the generality of our results. (4) There is still room for future improvements to further close the gap between upper and lower bounds. We believe this work opens several avenues for future research, including establishing the concrete lower bounds of risk-aware RLHF, improving the computational complexity of the algorithm, and conducting experiments in more diverse and interesting environments/simulations.
@@ Appendix A Notation
$\tilde{V}_{h}^{\pi}(s)$ Value function of state $s$ at step $h$ under policy $\pi$ in RLHF-MDP.
@@ Appendix B Risk Aware Object Computability
Unlike standard RL where each step’s reward can be observed, RLHF represents a type of RL characterized by once-per-episode feedback. As a result, our observable and estimable parameters are confined to the trajectory reward $r^{\star}_{\xi}$ and transition probability functions $\textbf{P}^{\star}$. Consequently, the traditional risk-aware objective might be unsuitable, as the reduction in available information prevents the computation of the original risk-aware measure. The risk measure selected for RLHF must satisfy the following condition: it should remain unique across MDPs where policies, trajectory rewards and transition probability functions even when each trajectory is fixed, but different step rewards, $r^{\star}(s,a)$, vary. This requirement ensures that the risk measure consistently reflects the underlying preferences regardless of variations in specific step rewards.
@@ B.1 Nested Object
Initial case. Using the tree-like RLHF-MDP algorithm and the initial conditions of the Bellman equation, at the final step $h=H$, we have
@@ E.1 Regret Lower Bound of Nested Reward
According to the definition of nested-CVaR-RLHF risk objective in Eq. 4, we have:
@@
According to the definition of nested-CVaR-RLHF risk objective in 4, we have:
@@
Appendix F The optimal policy calculation for known RLHF MDP
In this section, we describe how to compute the optimal policy for the CVaR objective when the RLHF-MDP is known; this approach is described in detail in Bäuerle & Ott (2011); Bastani et al. (2022). Following this work, we consider the setting where we are trying to minimize cost rather than maximize reward. In particular, consider an know RLHF MDP $\mathcal{M}(\mathcal{S},\mathcal{A},r_{\xi}^{\star},\mathbf{P}^{\star},H)$, where $\mathcal{S}$ and $\mathcal{A}$, and our goal is to compute a policy $\pi$ that maximizes its CVaR objective.
@@ F.1 Static CVaR object
Value iteration. we reconstruct the MDP as enlarge the state space $\tilde{s}_{h}=(\xi_{h},\rho)$, where $\rho$ will work as a quantile value. Letting $S_{1}$ be the initial state of the original RLHF MDP $\mathcal{M}$
@@ F.2 Nested CVaR object
Value iteration. we reconstruct the MDP as enlarge the state space $\tilde{s}_{h}=(\xi_{h},\rho)$, where $\rho$ will work as a quantile value. Letting $S_{1}$ be the initial state of the original RLHF MDP $\mathcal{M}$
@@ 3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contribution and scope: RA-RLHF algorithm.